A person invests x dollars at 5% APR (annual percentage rate) and y dollars at 4% APR. If the total annual income is $2000, solve for y as a function of x
Answer by mananth(16946)

5%x+4%y=2000
multiply by 100
5x+4y=200000
4y=200000-5x
y= 50000-(5/4)x